\begin{thebibliography}{99}
\bibitem[BL90]{BL}
M. Ben-Or and N. Linial. Collective Coin Flipping. In Randomness and Computation S. Micali ed.
Academic press, new York, 1990.
\bibitem[Tal94]{Tal94}
Talagrand, Michel. ``On Russo's approximate zero-one law." The Annals of Probability (1994): 1576-1587.
\bibitem[KKL88]{KKL88} Kahn, Jeff, Gil Kalai, and Nathan Linial. ``The influence of variables on Boolean functions." Foundations of Computer Science, 1988., 29th Annual Symposium on. IEEE, 1988.
\bibitem[BKK+92]{BKK+92}Bourgain, Jean, et al. ``The influence of variables in product spaces." Israel Journal of Mathematics 77.1-2 (1992): 55-64.
\bibitem[Fri98]{Fri98} Friedgut, Ehud. ``Influences in product spaces: KKL and BKKKL revisited." Combinatorics, Probability and Computing 13.01 (2004): 17-29.
\bibitem[DS05]{DS05} Dinur, Irit, and Samuel Safra. ``On the hardness of approximating minimum vertex cover." Annals of Mathematics (2005): 439-485.
\bibitem[CKK+06]{CKK+06} Chawla, Shuchi, et al. ``On the hardness of approximating multicut and sparsest-cut." computational complexity 15.2 (2006): 94-114.
\bibitem[KR08]{KR08} Khot, Subhash, and Oded Regev. ``Vertex cover might be hard to approximate to within $2- \epsilon$." Journal of Computer and System Sciences 74.3 (2008): 335-349.
\bibitem[FK96]{FK96} Friedgut, Ehud, and Gil Kalai. ``Every monotone graph property has a sharp threshold." Proceedings of the American mathematical Society 124.10 (1996): 2993-3002.
\bibitem[Fri99]{Fri99} Friedgut, Ehud ``Sharp thresholds of graph properties, and the $k$-sat problem." Journal of the American Mathematical Society 12.4 (1999): 1017-1054.
\bibitem[OW09]{OW09} Ryan O’Donnell and Karl Wimmer. KKL, Kruskal-Katona, and monotone nets. In Proceedings of the 50th
Annual IEEE Symposium on Foundations of Computer Science, pages 725–734, 2009.
\bibitem[KR09]{KR09}Krauthgamer, Robert, and Yuval Rabani. ``Improved lower bounds for embeddings into $l_1$." SIAM Journal on Computing 38.6 (2009): 2487-2498.
\bibitem[DKSV06]{DKSV06} Nikhil Devanur, Subhash Khot, Rishi Saket, and Nisheeth Vishnoi. Integrality gaps for Sparsest Cut and
Minimum Linear Arrangement problems. In Proceedings of the 38th Annual ACM Symposium on Theory
of Computing, pages 537–546, 2006.
\bibitem[OS08]{OS08} Ryan O’Donnell and Rocco Servedio. The Chow Parameters problem. In Proceedings of the 40th Annual
ACM Symposium on Theory of Computing, pages 517–526, 2008.
\bibitem[Raz95]{Raz95} Ran Raz. A parallel repetition theorem. In Proceedings of the 27th Annual ACM Symposium on Theory
of Computing, pages 447–456, 1995.
\end{thebibliography}

\documentclass[11pt]{article}

\usepackage{amsmath, amsthm, amsopn, amssymb, enumerate, ifpdf}
\usepackage[english]{babel}
\ifpdf
\usepackage[pdftex,urlcolor=blue]{hyperref}
\else
\usepackage[hypertex,urlcolor=blue]{hyperref}
\fi
\hypersetup{colorlinks}
\usepackage{geometry}
\newtheorem{thm}{Theorem}[]

\newtheorem{claim}[thm]{Claim}
\theoremstyle{definition}

\def\FullBox{\hbox{\vrule width 8pt height 8pt depth 0pt}}
\newcommand{\QED}{\;\;\;\FullBox}
\newenvironment{Proof}{\noindent{\bf Proof:~~}}{\hfill\QED}
\newcommand{\BPF}{\begin{Proof}} \newcommand {\EPF}{\end{Proof}}

\newcommand{\E}{\mathbb{E}}
\newcommand{\Var}{\text{Var}}
\renewcommand{\Pr}{{\rm Pr}}

\newcommand{\eqdef}{\stackrel{\rm def}{=}}

\DeclareMathOperator\Inf{Inf}
\renewcommand\emptyset{\phi}

\title{On the Converse of Talagrand's Influence Inequality}

\author{Saleet Klein\thanks{Tel Aviv University, \href{mailto:saleetklein@mail.tau.ac.il}{\nolinkurl{saleetklein@mail.tau.ac.il}}} \and Amit Levi\thanks{Tel Aviv University, \href{mailto:amitlev3@post.tau.ac.il}{\nolinkurl{amitlev3@post.tau.ac.il}}} \and Muli Safra\thanks{Tel Aviv University, \href{mailto:safra@cs.tau.ac.il}{\nolinkurl{safra@cs.tau.ac.il }}} \and Clara Shikhelman\thanks{Tel Aviv University, \href{mailto:clarashk@mail.tau.ac.il‬}{\nolinkurl{clarashk@mail.tau.ac.il‬}}} \and Yinon Spinka\thanks{Tel Aviv University, \href{mailto:yinonspi@post.tau.ac.il }{\nolinkurl{yinonspi@post.tau.ac.il}}} \\}

\date{}

\begin{document}
\maketitle
\begin{abstract}
In~\cite{Tal94}, Talagrand showed a generalization of the celebrated KKL theorem. In this work, we prove that the converse of this generalization also holds. Namely, for any sequence of numbers $0<a_1,a_2,\ldots,a_n\le 1$ such that $\sum_{j=1}^n a_j/(1-\log a_j)\ge C$ for some constant $C>0$, it is possible to find a roughly balanced Boolean function $f$ such that $\Inf_j[f] < a_j$ for every $1 \le j \le n$.
\end{abstract}
\section{Introduction}

In their seminal paper \cite{KKL88}, Kahn, Kalai and Linial showed that for any Boolean function $f$ there exists a coordinate $1 \le j \le n$ such that $\Inf_j[f]\ge c\cdot\frac{\log n}{n}\cdot \Var[f]$, where $c>0$ is a universal constant. This result, followed by the generalizations of Bourgain el al.\ \cite{BKK+92}, Talagrand \cite{Tal94} and Friedgut \cite{Fri98}, was a milestone for numerous results in different areas in computer science and mathematics such as hardness of approximation \cite{DS05,CKK+06,KR08}, distributed computing \cite{BL}, communication complexity \cite{Raz95}, metric embeddings \cite{KR09,DKSV06}, learning theory \cite{OS08,OW09}, random $k$-SAT \cite{Fri99}, random graphs \cite{FK96} and extremal combinatorics \cite{OW09}.

Talagrand's paper ``On Russo's approximate zero-one law" \cite{Tal94}, generalized KKL's result and stated that for every Boolean function $f$,
\[ \sum_{j=1}^n \frac{\Inf_j[f]}{1-\log\Inf_j[f]}\ge K\cdot \Var[f] ,\]
where $K>0$ is a universal constant.
We refer to this sum as \emph{Talagrand sum}.

We study whether the converse Talagrand's theorem holds. Namely, given a sequence of numbers, $0 < a_1,a_2,\ldots,a_n \le 1$ whose Talagrand sum is greater than a constant $C>0$, can one find a roughly balanced Boolean function $f$ such that $\Inf_j[f] < a_j$ for all $1 \le j \le n$. We show that this is true (up to a constant) not only for balanced functions, but also for unbalanced functions. 
\section{Main Result}
Let $n$ be a positive integer (which is henceforth fixed).
A {\em Boolean function} on $n$ variables is a function $f \colon \{0,1\}^n \to \{0,1\}$.
Let $x$ be a uniformly chosen element in $\{0,1\}^n$.
For $1 \le j \le n$, denote by $x^j$ the vector $x$ with the $j$-th coordinate flipped, i.e.,
\[ x^j \eqdef (x_1, \dots, x_{j-1}, 1-x_j, x_{j+1}, \dots, x_n) .\]
We define the {\em influence} of the $j$-th variable on $f$ to be
\[ \Inf_j[f] \eqdef \Pr[f(x) \neq f(x^j)] .\]
That is, $\Inf_j[f]$ is the probability that flipping the $j$-th bit affects the outcome of the function. 
We say that a Boolean function $f$ is {\em balanced} if $\Pr[f(x)=0]=1/2$. Throughout this paper, $\ln$ denotes the natural logarithm and $\log$ denotes the base-$2$ logarithm.

\begin{thm}
Let $0 < a_1,a_2,\ldots,a_n \le 1$ and denote
\[ \alpha \eqdef \sum_{j=1}^n\frac{a_j}{1-\log a_j} - 2 \ln 2 .\]
Then, for any $0<\mu \le 1-\exp(-\alpha/8)$, there exists a Boolean function $f$ on $n$ variables such that
\[ \mu \leq \E[f(x)] \le \frac{3}{4} \mu + \frac{1}{4} \]
and
\[ \Inf_j[f] < a_j \quad\text{ for all } 1 \le j \le n.\]
\end{thm}
\BPF
Following Ben-Or and Linial's example of tribes \cite{BL}, which minimizes the influence of each variable, we wish to construct a function in a similar manner by aggregating variables $x_j$ into tribes of various sizes according to the desired bound $a_j$ on their influence.

We may assume without loss of generality that $a_1 \ge a_2 \ge \cdots \ge a_n$. Define an integer $m \geq 0$ and a sequence of integers $k_0,k_1,\dots,k_m,k_{m+1}$ by
\begin{align*}
k_0 &\eqdef 0, \\
k_i &\eqdef \min \{ k \ge 1 ~:~ a_{k_0+\cdots+k_{i-1}+k} > 2^{1-k} \} , \quad 1 \leq i \leq m , \\
k_{m+1} &\eqdef n - (k_0 + \cdots + k_m) + 1 ,
\end{align*}
where $m$ is defined by the condition $\{ k \ge 1 ~:~ a_{k_0+\cdots+k_m+k} > 2^{1-k} \} = \emptyset$. Note that $2 \leq k_1 \le k_2 \le \cdots \le k_m$.
The integers $k_1,\dots,k_m$ represent the sizes of the tribes.
We shall show that there exists an integer $0 \leq m_* \leq m$ such that the function $f \colon \{0,1\}^n \to \{0,1\}$ defined by
\begin{equation}\label{eq:def-f}
f(x_1,\dots,x_n) \eqdef \bigvee_{i=1}^{m_*} \bigwedge_{j=1}^{k_i} x_{k_1+\cdots+k_{i-1}+j}
\end{equation}
satisfies the conclusion of the theorem.
We first show the following.
\begin{claim} \label{clm_expectation}
\[ \sum_{i=1}^{m} 2^{-k_i} \ge \frac{\alpha}{8} . \]
\end{claim}

\BPF
For $0 \leq i \leq m$, denote $s_i \eqdef k_0+\cdots+k_i$.
Note that, by the definition of $k_i$, we have
\[ a_{s_i + j} \le 2^{1-j} , \quad 0 \le i \le m,~ 1 \leq j \leq k_{i+1} - 1 \]
and, in particular,
\[ a_{s_i} \le a_{s_i - 1} = a_{s_{i-1}+k_i-1} \le 2^{2-k_i}, \quad 1 \le i \le m . \]
Therefore, since $x/(1- \log x)$ is increasing on $[0,1]$,
\begin{align*}
\sum_{j=1}^n\frac{a_j}{1-\log a_j}
 &= \sum_{j=1}^{k_1-1} \frac{a_j}{1-\log a_j}
  + \sum_{i=1}^{m} \sum_{j=0}^{k_{i+1}-1} \frac{a_{s_i+j}}{1-\log a_{s_i+j}} \\
 &\le \sum_{j=1}^{k_1-1} \frac{a_j}{1-\log a_j}
  + \sum_{i=1}^{m} \left( \sum_{j=0}^{k_i-1} \frac{a_{s_i}}{1-\log a_{s_i}} + \sum_{j=k_i}^{k_{i+1}-1} \frac{a_{s_i+j}}{1-\log a_{s_i+j}} \right) \\  
  &\le \sum_{j=1}^{k_1-1} \frac{2^{1-j}}{j}
  + \sum_{i=1}^{m} \left( k_i \frac{2^{2-k_i}}{k_i-1} + \sum_{j=k_i}^{k_{i+1}-1} \frac{2^{1-j}}{j} \right) \\  
  &= \sum_{j=1}^n \frac{2^{1-j}}{j}
  + 4 \sum_{i=1}^{m} \frac{k_i}{k_i-1} 2^{-k_i} \\
  &\le 2\ln 2 + 8 \sum_{i=1}^{m} 2^{-k_i} .
\end{align*}
Thus, the claim follows by the definition of $\alpha$.
\EPF

\medbreak

We construct the function as described in \eqref{eq:def-f}, where we keep adding tribes until we reach a point at which the expected value becomes at least $\mu$. We denote this tribe by $m_*+1$, so that at the end of the process we have $m_*$ tribes. Precisely, $m_*$ is defined by
\[ m_* \eqdef \min\left\{ 1 \le r \le m ~:~ \prod_{i=1}^r (1-2^{-k_i}) \leq 1-\mu \right\} .\]
Indeed, this is well-defined, since by Claim~\ref{clm_expectation},
\[ \prod_{i=1}^m (1-2^{-k_i}) \le \exp\left(- \sum_{i=1}^m 2^{-k_i} \right) \le \exp\left(-\frac{\alpha}{8}\right) \le 1-\mu .\]
Thus, with this choice of $m_*$, we have
\[ \E[f(x)] = \Pr[f(x)=1] = 1 - \Pr[f(x)=0] = 1 - \prod_{i=1}^{m_*} (1-2^{-k_i}) \ge \mu .\]
%
%
We now show that the expectation of the function $f$ constructed above is less than $\frac{3}{4} \mu + \frac{1}{4}$.
Indeed, by the minimality of $m_*$, and since $k_i \geq 2$ for all $1 \leq i \leq m$, we have
\[ \E[f(x)] = 1 - \prod_{i=1}^{m_*} (1-2^{-k_i}) < 1 - (1-\mu) (1-2^{-k_{m_*}}) \le 1 - \frac{3}{4}(1-\mu) = \frac{3}{4} \mu + \frac{1}{4} .\]

It remains to check that the influence of the $j$-th variable on $f$ is bounded above by $a_j$.
Indeed, if $s_{i-1} < j \leq s_i$ for some $1 \le i \le m_*$, then
\[ \Inf_j[f] = 2^{1-k_i} \prod_{\substack{\ell=1\\\ell \neq i}}^{m_*} (1-2^{-k_\ell}) \le 2^{1-k_i} < a_{s_i} \leq a_j .\]
Since $\Inf_j[f] = 0$ for all $s_{m_*} < j \le n$, the proof is complete.
\EPF

\section{Acknowledgments }
The authors would like to thank Gil Kalai for pointing out this question and to Guy Kindler for useful discussions.

\end{document}